OpenStudy (anonymous):

2 AND I THIRD + 7 OVER 2?? SHOW WORK

OpenStudy (anonymous):

7/3+7/2 14/6+21/6 35/6 5 and 5/6

OpenStudy (anonymous):

ITS 2 AND 1 THIRD AND 7 OVER 2!!

OpenStudy (anonymous):

2 and 1/3 1=3/3, 2=6/3, 6/3+1/3=7/3

OpenStudy (anonymous):

7/3=2 1/3

OpenStudy (anonymous):

WORK IT OUT AGAIN FOR ME PLEASE!!!

OpenStudy (anonymous):

okay 2 1/3 + 7/2 7/3 + 7/2 (change 2 1/3 to a fraction so it is easier to add) 14/6 + 21/6 (cross multiply to get the same denominator so you can add them) 35/6 (add the numerators) 5 5/6 (simplify 6*5=30 with a remainder of 5)